# A Diffusion-Based Approach for Simulating Forward-in-Time State-Dependent Speciation and Extinction Dynamics

### v1

@@ 2.5 Comparison on diffusion-based and tree-based models using simulation
In this section we show that our diffusion-based approach correctly models the temporal behaviour of range state frequencies in a GeoSSE model. To validate, we compare our results with a tree-based approach that explicitly simulates phylogenetic trees under the same GeoSSE parameter values using the MASTER package [38] implemented in BEAST2 [2]. Details for setting up reaction equations for the MASTER simulation can be found in Appendix 8.5.
@@
    where $dW_{t}$ is a Wiener process. We draw a sample path by using the following approximation,
@@
    Compute the total number of species at $t+\Delta t\in[0,T]$
@@
    Given the output from the previous steps, and the following Itô SDE with the initial frequency of species of range state $i$, $\Pi_{i}(0)=\frac{N_{i}(0)}{N(0)}$,
@@
    where $dW_{t}$ is a Wiener process. We draw a sample path by using the following approximation,
@@
    where $\sqrt{\Delta t}U_{t}\sim\sqrt{\Delta t}N(0,1)$ is a (discretized) standard Wiener process; $\mu_{\Pi_{i}}(t)$ and $\sigma_{\Pi_{i}}(t)$ are computed using Eqs (10)-(11), respectively.
@@
    We consider a 3-region GeoSSE model, then we simulate range state dynamics using tree-based approach (via the MASTER package in BEAST2) and the diffusion-based approach over 1000 replicates on $[0,10]$ time interval. For diffusion-based approach, at each time step, we perform a rejection sampling for each range state if the count value for a state is less than 00 since the number of species with any range state cannot be negative. We consider the following scenarios for the GeoSSE model,
@@
    We visualize the trajectory of mean state counts for each range state from both diffusion and tree-based approaches. We also visualize stacked bar charts of expected state frequencies for both approaches. To compute the state frequencies under the tree-based approach across replicates, we use the following analytical formula
@@
    We simulate frequency trajectories under the diffusion-based approach using Eq. (49).
@@ 2.6 Deriving rate parameters that lead to stationary state frequencies when $N$ is large
Knowing that $\Pi_{i}=\frac{N_{i}}{N}$, we re-write Eqs. (26)-(37) as follows,
@@
Furthermore, it is biologically reasonable that all rate parameters must be positive, as all modeled events have some non-zero probability of occurring. That is,
@@
where $r_{W^{+}_{i}},r_{D^{+}_{i}},r_{B^{+}_{i}},r_{E^{+}_{i}},r_{D^{-}_{i}},r_{B^{-}_{i}},r_{E^{-}_{i}}$ consist of sums of rates across all adjacent states that correspond to the events $W^{+}_{i},D^{+}_{i},B^{+}_{i},E^{+}_{i},D^{-}_{i},B^{-}_{i},E^{-}_{i}$, respectively.
@@ 2.7 Deriving stationary state frequencies given rate parameters in a GeoSSE model
Consider a 2-region GeoSSE model with state space $S=\{\{A\},\{B\},\{A,B\}\}$. Given the rate parameters from the model and initial state frequencies, $\Pi_{A}(0)=\Pi_{A}^{0}$, $\Pi_{B}(0)=\Pi_{B}^{0}$, $\Pi_{AB}(0)=\Pi_{AB}^{0}$, the general solution to Eq. (79) is given by,
$\bm{\Pi}$ $=$ $\begin{bmatrix}\Pi_{A}(t)\\ \Pi_{B}(t)\end{bmatrix}$ (83)
$=$ $C_{1}\bm{\nu}_{1}e^{\lambda_{1}t}+C_{2}\bm{\nu}_{2}e^{\lambda_{2}t}+\bm{K},$ (84)
and $\Pi_{AB}(t)=1-\Pi_{A}(t)-\Pi_{B}(t)$, provided that $\Pi_{A}(t)+\Pi_{B}(t)\leq 1$.
@@
$\hat{\Pi}_{A}$ $=$ $\frac{num_{A}}{denom_{A}},$ (85)
$\hat{\Pi}_{B}$ $=$ $1-\left(\frac{e_{A}+d_{AB}+b^{A}_{B}+e_{B}}{w_{A}+b^{A}_{B}+e_{B}}\right)\left(\frac{num_{A}}{denom_{A}}\right),\hskip 20.00003pt$ (86)
$\hat{\Pi}_{AB}$ $=$ $1-\hat{\Pi}_{A}-\hat{\Pi}_{B},$ (87)
@@
$\bm{K}$ $=$ $\begin{bmatrix}\hat{\Pi}_{A}\\[10.0pt] \hat{\Pi}_{B}\end{bmatrix},$ (96)
@@ 2.8 Deriving time to reach stationary state frequencies in a GeoSSE model
In this section, we describe a method for deriving time to reach stationary state frequencies in a 2-region GeoSSE model.
@@ 3.1 Diffusion-based approach is a good approximation to tree-based approach for describing state dynamics
In this section, we visualize the range state dynamics using tree-based and diffusion-based approaches under several GeoSSE scenarios described in Section 2.5 (Figs. 3-6). In all these scenarios, we show that the null hypothesis that the average counts of the ranges states at the end of the simulation time between these approaches are equal cannot be rejected (Table 1). This shows that the diffusion-based approach is a good approximation for means to the tree-based approach.
@@
Figure 3: Top & middle panels: the trajectories of average count of range states for endemic species (Figs. (a)-(c)) and widespread species (Figs. (d)-(g)) over $[0,10]$ time interval and over 1000 simulations runs for the three-region GeoSSE model as described in Example 1 each simulated under both diffusion-based process (red line) and tree-based process (black line). The gray trajectories show the dynamics across 1000 replicates simulated under diffusion-based process. Bottom panel: stacked bar chart showing the state frequencies over time using diffusion-based approach (Fig. (g)) and tree-based approach (Fig. (h)). Simulations are conducted using the following parameter values: $w_{A}=w_{B}=w_{C}=0.03,b^{A}_{B}=0.08,b^{A}_{C}=0.10,b^{B}_{C}=0.06,b^{A}_{BC}=0.04,b^{B}_{AC}=0.12,b^{C}_{AB}=0.06,e_{A}=e_{B}=e_{C}=0,d_{AB}=d_{BA}=d_{AC}=d_{CA}=d_{BC}=d_{CB}=0$.
@@
Figure 4: Top & middle panels: the trajectories of average count of range states for endemic species (Figs. (a)-(c)) and widespread species (Figs. (d)-(g)) over $[0,10]$ time interval and over 1000 simulations runs for the three-region GeoSSE model as described in Example 2 simulated under both diffusion-based process (red line) and tree-based process (black line). The gray trajectories show the dynamics across 1000 replicates simulated under diffusion-based process. Bottom panel: stacked bar chart showing the state frequencies over time using diffusion-based approach (Fig. (g)) and tree-based approach (Fig. (h)). Simulations are conducted using the following parameter values: $w_{A}=w_{B}=w_{C}=0.03,b^{A}_{B}=b^{A}_{C}=b^{B}_{C}=b^{A}_{BC}=b^{B}_{AC}=b^{C}_{AB}=0,e_{A}=e_{B}=e_{C}=0,d_{AB}=d_{BA}=0.03,d_{AC}=d_{CA}=0.04,d_{BC}=d_{CB}=0.05$.
@@
Figure 5: Top & middle panels: the trajectories of average count of range states for endemic species (Figs. (a)-(c)) and widespread species (Figs. (d)-(g)) over $[0,10]$ time interval and over 1000 simulations runs for the three-region GeoSSE model as described in Example 3 simulated under both diffusion-based process (red line) and tree-based process (black line). The gray trajectories show the dynamics across 1000 replicates simulated under diffusion-based process. Bottom panel: stacked bar chart showing the state frequencies over time using diffusion-based approach (Fig. (g)) and tree-based approach (Fig. (h)). Simulations are conducted using the following parameter values: $w_{A}=w_{B}=w_{C}=0.03,b^{A}_{B}=b^{A}_{C}=b^{B}_{C}=b^{A}_{BC}=b^{B}_{AC}=b^{C}_{AB}=0,e_{A}=0.01,e_{B}=0.02,e_{C}=0.025,d_{AB}=d_{BA}=d_{AC}=d_{CA}=d_{BC}=d_{CB}=0$.
@@
Figure 6: Top & middle panels: the trajectories of average count of range states for endemic species (Figs. (a)-(c)) and widespread species (Figs. (d)-(g)) over $[0,10]$ time interval and over 1000 simulations runs for the three-region GeoSSE model as described in Example 4 simulated under the diffusion-based process (red line) and tree-based process (black line). The gray trajectories show the dynamics across 1000 replicates simulated under diffusion-based process. Bottom panel: stacked bar chart showing the state frequencies over time using diffusion-based approach (Fig. (g)) and tree-based approach (Fig. (h)). Simulations are conducted using the following parameter values: $w_{A}=0.09,w_{B}=0.06,w_{C}=0.07,b^{A}_{B}=b^{A}_{C}=b^{B}_{C}=b^{A}_{BC}=b^{B}_{AC}=b^{C}_{AB}=0.04,e_{A}=0.002,e_{B}=0.003,e_{C}=0.001,d_{AB}=d_{BA}=0.006,d_{AC}=d_{CA}=0.003,d_{BC}=d_{CB}=0.001$.
@@
Table 1: The sample mean count for each range state at the end of simulation time, $\bar{N}_{i,end}$, computed under tree-based and diffusion-based simulations across different GeoSSE scenarios described in Section 2.5. The “Lower bound” and “Upper bound” represent the $95\%$ confidence interval of the average count for each range state using diffusion and tree based approaches. The “$95\%$ CI $r_{i,var}$” correspond to the $95\%$ confidence interval of the ratio of two sample variances from diffusion and tree based approaches for range state $i$. $p_{var}$ and $p_{mean}$ correspond to $p$ value from the $F$ test and the Welch’s unequal variances t-test, respectively.

### v2

@@ 2.5 Comparison on diffusion-based and tree-based models using simulation
In this section we show that our diffusion-based approach correctly models the temporal behaviour of range state frequencies in a GeoSSE model. To validate, we compare our results with a tree-based approach that explicitly simulates phylogenetic trees under the same GeoSSE parameter values using the MASTER package [40] implemented in BEAST2 [2]. When simulating given a large number of species initially, $N(0)>>0$, both diffusion-based and MASTER-based simulations are conditioned only for the process to run until a specific elapsed time $T$. Later in Appendix 7.6, when we simulate using both approaches starting with a single species in random state, $N(0)=1$, we condition the process under both elapsed time and survival until the present. Details for setting up reaction equations for the MASTER simulation can be found in Appendix 7.5.
@@
    where $dW_{t}$ is a Wiener process, we draw a sample path by using the following approximation,
@@
    Given $N_{i}(t+\Delta t)$ for each $i\in S$ from step $1$, we compute the total number of species at $t+\Delta t\in[0,T]$
@@
    Next, using $N_{i}(t)$ and $N(t)$ from steps 1-2, we compute the infinitesimal mean, $\mu_{\Pi_{i}}(t)$, and infinitesimal variance, $\sigma_{\Pi_{i}}(t)$ using Eqs. (9)-(10), respectively. Given $\mu_{\Pi_{i}}(t)$, $\sigma_{\Pi_{i}}(t)$, and the following Itô SDE with the initial frequency of species of range state $i$, $\Pi_{i}(0)=\frac{N_{i}(0)}{N(0)}$,
@@
    where $dW_{t}$ is a Wiener process, we draw a sample path by using the following approximation,
@@
    where $\sqrt{\Delta t}U_{t}\sim\sqrt{\Delta t}N(0,1)$ is a (discretized) standard Wiener process.
@@
    We consider a 3-region GeoSSE model, then we simulate range state dynamics using tree-based approach (via the MASTER package in BEAST2) and the diffusion-based approach over 1000 replicates on $[0,10]$ time interval with 1000 time steps. Note that if one simulates over a longer time interval, then one needs to choose larger time steps to reduce the chance that multiple events occur within $\Delta t$ for the diffusion-based approach. For diffusion-based approach, at each time step, we assign a zero value to any state with a count less than zero since the number of species in any range states cannot be negative. This is reasonable because if $N_{i}(t)=0$, then some events are not permitted such as a local extinction. Note that although some $N_{i}$’s might be equal to 00, it is very unlikely for the whole clade to become extinct, i.e., $N(t)=0$, given a relatively large clade size at the beginning of each process (Fig. 4) and value of each parameter we pick for the simulations (Figs. 5-8). We consider the following scenarios for the GeoSSE model,
@@
    Figure 4: For each diffusion-based and MASTER-based simulation, we assume that we start each $N_{i}(t)$ simulation, given a relatively large clade size at the beginning, $N(0)>>0$
@@
    We visualize the trajectory of mean state counts for each range state from both diffusion and tree-based approaches. For each simulation, we start the forward-in-time simulation given relatively large clade size for diffusion-based approach to accurately predict the dynamics given by tree-based approach from MASTER simulations. We also visualize stacked bar charts of expected state frequencies for both approaches. To compute the state frequencies under the tree-based approach across replicates, we use the following analytical formula
@@
    We simulate frequency trajectories under the diffusion-based approach using Eq. (19). Also for diffusion-based approach, we normalize $\Pi_{i}(t)$ at each time step for each $i\in S$. Thus, keeping $\Pi_{i}(t)\leq 1$ at any time.
@@
    While all the diffusion-based simulations presented in the main text assume that we always start with a relatively large clade size, this is not how phylogenetic trees are normally simulated. Instead, most simulations generate the entire clade, beginning with one stem or two sister lineages to represent the origin of the process. However, the diffusion approximation assumes the number of species is large. Therefore, to adapt our diffusion-based model for clade-generation scenarios where the initial number of species is small, we adapted our diffusion-based simulation method to start the process with a single species in a random state (see Appendix 7.6). We show that the difference between diffusion-based and tree-based simulations is reduced after applying the correction.
@@ 2.6 Deriving rate parameters that lead to stationary state frequencies when $N$ is large
Knowing that $\Pi_{i}=\frac{N_{i}}{N}$, we re-write Eqs. (2.4)-(14) as follows,
@@
Furthermore, we assume all rate parameters must be positive, as all modeled events have some non-zero probability of occurring. That is,
@@
where $r_{W^{+}_{i}},r_{D^{+}_{i}},r_{B^{+}_{i}},r_{E^{+}_{i}},r_{D^{-}_{i}},r_{B^{-}_{i}},r_{E^{-}_{i}}$ consist of sums of rates across all adjacent states that correspond to the events $W^{+}_{i},D^{+}_{i},B^{+}_{i},E^{+}_{i},D^{-}_{i},B^{-}_{i},E^{-}_{i}$, respectively. $\Phi_{total,i}$ can also be thought as a flux for range state $i$. That is, it is a difference between total incoming rates and outgoing rates. For example, in a two-region GeoSSE model, we can define $\Phi_{total,\{A\}}$ as follows,
$\Phi_{total,\{A\}}=\left(2w_{A}+0+b^{A}_{B}+e_{B}\right)-(d_{AB}+0+e_{A}),$
where we have $r_{W^{+}_{\{A\}}}=2w_{A}$ because within-region speciation rate $w_{A}$ is acting on both endemic species with state $\{A\}$ and widespread species with state $\{A,B\}$.
@@ 2.7 Deriving stationary state frequencies given rate parameters in a GeoSSE model
Consider a 2-region GeoSSE model with state space $S=\{\{A\},\{B\},\{A,B\}\}$. Given the rate parameters from the model and initial state frequencies, $\Pi_{\{A\}}(0)=\Pi_{A}^{0}$, $\Pi_{\{B\}}(0)=\Pi_{B}^{0}$, $\Pi_{\{A,B\}}(0)=\Pi_{AB}^{0}$, the general solution to Eq. (26) is given by,
$\bm{\Pi}$ $=$ $\begin{bmatrix}\Pi_{\{A\}}(t)\\ \Pi_{\{B\}}(t)\end{bmatrix}$ (28)
$=$ $C_{1}\bm{\nu}_{1}e^{\lambda_{1}t}+C_{2}\bm{\nu}_{2}e^{\lambda_{2}t}+\bm{K},$
and $\Pi_{\{A,B\}}(t)=1-\Pi_{\{A\}}(t)-\Pi_{\{B\}}(t)$, provided that $\Pi_{\{A\}}(t)+\Pi_{\{B\}}(t)\leq 1$.
@@
$\hat{\Pi}_{\{A\}}$ $=$ $\frac{num_{A}}{denom_{A}},$ (29)
$\hat{\Pi}_{\{B\}}$ $=$ $1-\left(\frac{e_{A}+d_{AB}+b^{A}_{B}+e_{B}}{w_{A}+b^{A}_{B}+e_{B}}\right)\left(\frac{num_{A}}{denom_{A}}\right),\hskip 20.00003pt$ (30)
$\hat{\Pi}_{\{A,B\}}$ $=$ $1-\hat{\Pi}_{\{A\}}-\hat{\Pi}_{\{B\}},$ (31)
@@
$\bm{\nu}_{1}$ $=$ $\begin{bmatrix}-\frac{1}{2\left(b^{A}_{B}+e_{A}+w_{B}\right)}\left(-d_{AB}+d_{BA}-R\right)\\[10.0pt] 1\end{bmatrix},$
$\bm{\nu}_{2}$ $=$ $\begin{bmatrix}-\frac{1}{2\left(b^{A}_{B}+e_{A}+w_{B}\right)}\left(-d_{AB}+d_{BA}+R\right)\\[10.0pt] 1\end{bmatrix},$
$\bm{K}$ $=$ $\begin{bmatrix}\hat{\Pi}_{\{A\}}\\[10.0pt] \hat{\Pi}_{\{B\}}\end{bmatrix},$
@@ 2.8 Deriving time to reach stationary state frequencies in a GeoSSE model
In this section, we describe a method for deriving time to reach stationary state frequencies in a 2-region GeoSSE model. Note that we have assumed a relatively large clade size at the start of the process for simulating $\Pi_{i}(t)$. Thus, the following is time to stationary frequencies since some relatively large clade size (Fig. 4).
@@ 3.1 Diffusion-based approach is a good approximation to tree-based approach for describing state dynamics
In this section, we visualize the range state dynamics using tree-based and diffusion-based approaches under several GeoSSE scenarios described in Section 2.5 (Figs. 5-8). In all these scenarios, we show that the null hypothesis that the average counts of the ranges states at the end of the simulation time between these approaches are equal cannot be rejected (Table 1). This shows that the diffusion-based approach is a good approximation for means to the tree-based approach.
@@
Figure 5: Top & middle panels: the trajectories of average count of range states for endemic species (Figs. 5(a)-(c)) and widespread species (Figs. 5(d)-(g)) over $[0,10]$ time interval and over 1000 simulations runs for the three-region GeoSSE model as described in Example 1 each simulated under both diffusion-based process (red line) and tree-based process (black line). The gray trajectories show the dynamics across 1000 replicates simulated under diffusion-based process. Bottom panel: stacked bar chart showing the state frequencies over time using diffusion-based approach (Fig. 5(h)) and tree-based approach (Fig. 5(i)). In both approaches, we start the process with $N(0)=40$ and the following initial state frequencies: $\Pi_{\{A\}}(0)=\Pi_{\{B\}}(0)=\Pi_{\{C\}}(0)=\Pi_{\{A,B\}}(0)=\Pi_{\{A,C\}}(0)=\Pi_{\{B,C\}}(0)=\Pi_{\{A,B,C\}}(0)=\frac{1}{7}$. At $t=10$, the mean frequencies for each range state from both diffusion-based and tree-based simulations are as follows:
$\bar{\Pi}^{diffusion}_{\{A\}}=0.29$, $\bar{\Pi}^{tree}_{\{A\}}=0.29$; $\bar{\Pi}^{diffusion}_{\{B\}}=0.29$, $\bar{\Pi}^{tree}_{\{B\}}=0.29$;
$\bar{\Pi}^{diffusion}_{\{C\}}=0.29$, $\bar{\Pi}^{tree}_{\{C\}}=0.28$; $\bar{\Pi}^{diffusion}_{\{A,B\}}=0.04$, $\bar{\Pi}^{tree}_{\{A,B\}}=0.04$;
$\bar{\Pi}^{diffusion}_{\{A,C\}}=0.04$, $\bar{\Pi}^{tree}_{\{A,C\}}=0.04$; $\bar{\Pi}^{diffusion}_{\{B,C\}}=0.04$, $\bar{\Pi}^{tree}_{\{B,C\}}=0.05$;
$\bar{\Pi}^{diffusion}_{\{A,B,C\}}=0.01$, $\bar{\Pi}^{tree}_{\{A,B,C\}}=0.01$.
Simulations are conducted using the following parameter values: $w_{A}=w_{B}=w_{C}=0.03,b^{A}_{B}=0.08,b^{A}_{C}=0.10,b^{B}_{C}=0.06,b^{A}_{BC}=0.04,b^{B}_{AC}=0.12,b^{C}_{AB}=0.06,e_{A}=e_{B}=e_{C}=0,d_{AB}=d_{BA}=d_{AC}=d_{CA}=d_{BC}=d_{CB}=0$
@@
Figure 6: Top & middle panels: the trajectories of average count of range states for endemic species (Figs. 6(a)-(c)) and widespread species (Figs. 6(d)-(g)) over $[0,10]$ time interval and over 1000 simulations runs for the three-region GeoSSE model as described in Example 2 simulated under both diffusion-based process (red line) and tree-based process (black line). The gray trajectories show the dynamics across 1000 replicates simulated under diffusion-based process. Bottom panel: stacked bar chart showing the state frequencies over time using diffusion-based approach (Fig. 6(h)) and tree-based approach (Fig. 6(i)). In both approaches, we start the process with $N(0)=40$ and the following initial state frequencies: $\Pi_{\{A\}}(0)=\Pi_{\{B\}}(0)=\Pi_{\{C\}}(0)=\Pi_{\{A,B\}}(0)=\Pi_{\{A,C\}}(0)=\Pi_{\{B,C\}}(0)=\frac{1}{6}$, $\Pi_{\{A,B,C\}}(0)=0$. At $t=10$, the mean frequencies for each range state from both diffusion-based and tree-based simulations are as follows:
$\bar{\Pi}^{diffusion}_{\{A\}}=0.14$, $\bar{\Pi}^{tree}_{\{A\}}=0.15$; $\bar{\Pi}^{diffusion}_{\{B\}}=0.14$, $\bar{\Pi}^{tree}_{\{B\}}=0.14$;
$\bar{\Pi}^{diffusion}_{\{C\}}=0.13$, $\bar{\Pi}^{tree}_{\{C\}}=0.13$; $\bar{\Pi}^{diffusion}_{\{A,B\}}=0.08$, $\bar{\Pi}^{tree}_{\{A,B\}}=0.09$;
$\bar{\Pi}^{diffusion}_{\{A,C\}}=0.11$, $\bar{\Pi}^{tree}_{\{A,C\}}=0.11$; $\bar{\Pi}^{diffusion}_{\{B,C\}}=0.13$, $\bar{\Pi}^{tree}_{\{B,C\}}=0.13$;
$\bar{\Pi}^{diffusion}_{\{A,B,C\}}=0.27$, $\bar{\Pi}^{tree}_{\{A,B,C\}}=0.25$.
Simulations are conducted using the following parameter values: $w_{A}=w_{B}=w_{C}=0.03,b^{A}_{B}=b^{A}_{C}=b^{B}_{C}=b^{A}_{BC}=b^{B}_{AC}=b^{C}_{AB}=0,e_{A}=e_{B}=e_{C}=0,d_{AB}=d_{BA}=0.03,d_{AC}=d_{CA}=0.04,d_{BC}=d_{CB}=0.05$
@@
Figure 7: Top & middle panels: the trajectories of average count of range states for endemic species (Figs. 7(a)-(c)) and widespread species (Figs. 7(d)-(g)) over $[0,10]$ time interval and over 1000 simulations runs for the three-region GeoSSE model as described in Example 3 simulated under both diffusion-based process (red line) and tree-based process (black line). The gray trajectories show the dynamics across 1000 replicates simulated under diffusion-based process. Bottom panel: stacked bar chart showing the state frequencies over time using diffusion-based approach (Fig. 7(h)) and tree-based approach (Fig. 7(i)). In both approaches, we start the process with $N(0)=40$ and the following initial state frequencies: $\Pi_{\{A\}}(0)=\Pi_{\{B\}}(0)=\Pi_{\{C\}}(0)=\Pi_{\{A,B\}}(0)=\Pi_{\{A,C\}}(0)=\Pi_{\{B,C\}}(0)=\Pi_{\{A,B,C\}}(0)=\frac{1}{7}$, the mean frequencies for each range state from both diffusion-based and tree-based simulations are as follows:
$\bar{\Pi}^{diffusion}_{\{A\}}=0.26$, $\bar{\Pi}^{tree}_{\{A\}}=0.26$; $\bar{\Pi}^{diffusion}_{\{B\}}=0.23$, $\bar{\Pi}^{tree}_{\{B\}}=0.23$;
$\bar{\Pi}^{diffusion}_{\{C\}}=0.21$, $\bar{\Pi}^{tree}_{\{C\}}=0.21$; $\bar{\Pi}^{diffusion}_{\{A,B\}}=0.09$, $\bar{\Pi}^{tree}_{\{A,B\}}=0.09$;
$\bar{\Pi}^{diffusion}_{\{A,C\}}=0.08$, $\bar{\Pi}^{tree}_{\{A,C\}}=0.08$; $\bar{\Pi}^{diffusion}_{\{B,C\}}=0.07$, $\bar{\Pi}^{tree}_{\{B,C\}}=0.07$;
$\bar{\Pi}^{diffusion}_{\{A,B,C\}}=0.06$, $\bar{\Pi}^{tree}_{\{A,B,C\}}=0.06$.
Simulations are conducted using the following parameter values: $w_{A}=w_{B}=w_{C}=0.03,b^{A}_{B}=b^{A}_{C}=b^{B}_{C}=b^{A}_{BC}=b^{B}_{AC}=b^{C}_{AB}=0,e_{A}=0.01,e_{B}=0.02,e_{C}=0.025,d_{AB}=d_{BA}=d_{AC}=d_{CA}=d_{BC}=d_{CB}=0$
@@
Figure 8: Top & middle panels: the trajectories of average count of range states for endemic species (Figs. 8(a)-(c)) and widespread species (Figs. 8(d)-(g)) over $[0,10]$ time interval and over 1000 simulations runs for the three-region GeoSSE model as described in Example 4 simulated under the diffusion-based process (red line) and tree-based process (black line). The gray trajectories show the dynamics across 1000 replicates simulated under diffusion-based process. Bottom panel: stacked bar chart showing the state frequencies over time using diffusion-based approach (Fig. 8(h)) and tree-based approach (Fig. 8(i)). In both approaches, we start the process with $N(0)=40$ and the following initial state frequencies: $\Pi_{\{A,B\}}(0)=\Pi_{\{A,C\}}(0)=\Pi_{\{B,C\}}(0)=\Pi_{\{A,B,C\}}(0)=\frac{1}{4}$ and $\Pi_{\{A\}}(0)=\Pi_{\{B\}}(0)=\Pi_{\{C\}}(0)=0$. At $t=10$, the mean frequencies for each range state from both diffusion-based and tree-based simulations are as follows:
$\bar{\Pi}^{diffusion}_{\{A\}}=0.33$, $\bar{\Pi}^{tree}_{\{A\}}=0.33$; $\bar{\Pi}^{diffusion}_{\{B\}}=0.21$, $\bar{\Pi}^{tree}_{\{B\}}=0.21$;
$\bar{\Pi}^{diffusion}_{\{C\}}=0.25$, $\bar{\Pi}^{tree}_{\{C\}}=0.25$; $\bar{\Pi}^{diffusion}_{\{A,B\}}=0.06$, $\bar{\Pi}^{tree}_{\{A,B\}}=0.06$;
$\bar{\Pi}^{diffusion}_{\{A,C\}}=0.06$, $\bar{\Pi}^{tree}_{\{A,C\}}=0.06$; $\bar{\Pi}^{diffusion}_{\{B,C\}}=0.05$, $\bar{\Pi}^{tree}_{\{B,C\}}=0.05$;
$\bar{\Pi}^{diffusion}_{\{A,B,C\}}=0.03$, $\bar{\Pi}^{tree}_{\{A,B,C\}}=0.03$.
Simulations are conducted using the following parameter values: $w_{A}=0.09,w_{B}=0.06,w_{C}=0.07,b^{A}_{B}=b^{A}_{C}=b^{B}_{C}=b^{A}_{BC}=b^{B}_{AC}=b^{C}_{AB}=0.04,e_{A}=0.002,e_{B}=0.003,e_{C}=0.001,d_{AB}=d_{BA}=0.006,d_{AC}=d_{CA}=0.003,d_{BC}=d_{CB}=0.001$
@@
Table 1: The sample mean count for each range state at the end of simulation time, $\bar{N}_{i,end}$, computed under tree-based and diffusion-based simulations across different GeoSSE scenarios described in Section 2.5. The “Lower bound” and “Upper bound” represent the $95\%$ confidence interval of the average count for each range state using diffusion and tree based approaches. The “$95\%$ CI $r_{i,var}$” correspond to the $95\%$ confidence interval of the ratio of two sample variances from diffusion and tree based approaches for range state $i$. $p_{var}$ and $p_{mean}$ correspond to $p$ value from the $F$ test and the Welch’s unequal variances t-test, respectively